• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name slifer

    Log entry time 17:07:10 on October 3,1998

    Entry number 2241

    keyword=target


    16:42 NMR measurement. Cell Temp (upper pumping )had dropped
    to 205 C. Waited for oven to stabilize.
    16:58 NMR - 10.2 mV (fitted) , ~33%/EPR